OpenStudy (anonymous):

Consider the parametric equation x = 3(cos(theta) + (theta)sin(theta) y = 3(sin(theta) - (theta)cos(theta) What is the length of the curve for theta= 0 to theta= (5/10)pi?

OpenStudy (amistre64):

ds when using polars is something like: sqrt(r^2 + (t')^2)

OpenStudy (amistre64):

or we might be able to use: sqrt(x'^2 + y'^2)

OpenStudy (amistre64):

t is easier to write instead of theta all the time

OpenStudy (amistre64):

x = 3(cos(t) + t sin(t) ) x' = 3(-sin(t)+sin(t)+tcos(t)) = 3t cos(t) y = 3(sin(t) - t cos(t) ) y' = 3(cos(t) -cos(t)+t sin(t) ) = 3t sin(t)

OpenStudy (amistre64):

x'^2 = 9t^2 cos^2(t) y'^2 = 9t^2 sin^2(t) add em up 9t^2(cos^2(t)+sin^2(t)) = 9t^2 sqrt(9t^2) = 3t integrate 3t from 0 to pi/2

OpenStudy (anonymous):

Got It! Thanks a lot!